# A Survey of Reinforcement Learning from Human Feedback

### v1

@@ 3 Feedback
A feedback type is a kind of interaction in which a human conveys some information about their preferences. Examples include pairwise comparisons and direct critiques. This section is concerned with the many ways in which human feedback can be expressed and used. Several previous works have already studied and sorted feedback types by listing the most common ones [125] and discussing their attributes and dimensions [181, 167]. The attributes and classes described in this section build upon this prior work and can be considered a synthesis and extension of it.
@@ 3.1 Attributes of Feedback Types
Feedback types may differ on many dimensions, some of which relate to the way feedback is given (arity, involvement), others to the form of the query instance it is given on (granularity, abstraction), and yet others to features of the human interaction (intent, explicitness). The attributes we discuss in this section are based on the framework proposed by [181]. We have adjusted and added terminology where it aids clarity, generalized the distinction between relative and absolute feedback to ‘arity’, and added the categories of co-generative involvement and literal intent. We have also systematically analyzed a set of exemplary feedback types with respect to these dimensions (see the next section), extending the initial work in that direction by [181] who consider a smaller set of abstract classes. Furthermore, we systematically analyze a set of exemplary feedback types in the next section, expanding on the initial examination of a smaller range of abstract classes by [181]. In the following, we will introduce each of the six attributes in more detail.
@@
Feedback may also differ on the granularity of the instances being evaluated. This ranges from whole episode recordings over partial segments to feedback on individual steps (i.e., states, actions, or state-action pairs). A more coarse-grained granularity has the advantage of giving humans more context and getting feedback for larger sections of behavior but also poses credit assignment problems. Finer-grained feedback is much easier to learn from and conveys more information, but is often impractical or tedious for humans to provide. Note that we only classify a type of feedback as “episode” granularity if it requires entire episodes. If it is compatible with partial segments as well, we classify it as “segment” even if the discussed source paper uses entire episodes.
@@
Intent
The assumed human intent can be important for feedback processing. A human may be evaluative, instructive, or descriptive in their explicit feedback, while they are generally literal in their implicit feedback. Evaluative, instructive, and descriptive feedback is pedagogical in nature, aiming to teach a reward function, whereas literal feedback is a byproduct of a human actor’s efforts to optimize the reward function directly. While evaluative, instructive, and literal feedback is generally given within the context of a particular query instance, descriptive feedback can describe the task more generally, e.g., through a partial reward function.
The distinction between literal and pedagogical feedback was introduced by [184]. They argue that humans with pedagogical intent (i.e., evaluative, instructive, or descriptive) may act differently compared to humans with literal intent. Even though they find that assuming the (wrong) literal intent can still lead to better reward inference, it still indicates that it can be important to know this intent to choose the right human model (see Section 5.1.5).
@@ 3.2 Common Classes
Table 4: An overview of the common classes and their defining attributes. When an attribute is not specified, this means that it is not a defining feature of the class and may vary in different instantiations.
Class Granularity Involvement Arity Abstraction Intent Explicitness
@@
Task Desc. Episode Observed Feature Descriptive Explicit
@@
3.2.1 Critique
@@
There are many examples of critique feedback in the literature. [277] employ binary feedback on individual state and action pairs. Although they learn a shaping reward that complements an environment reward signal, the same technique could be used without environment reward. [118] extend this to multi-label feedback, allowing the user to distinguish between a regular good or bad action and a terminal action that achieves the goal or fails to do so. They map these classes to scalar values and then learn a reward model by regression. [263] present an approach to learning a reward model from noisy critiques in the form of human physiological signals (brain signals) using active querying. In contrast to this action-level feedback, [81, 237] rely on binary outcome success labels. [81] introduce the basic approach, which [237] extend by moving to an off-policy setting and including online queries, thereby reducing the reliance on many positive examples by interactively correcting false positives.
@@
In addition to learning the main reward function, critique can also be used for safety evaluation. [60] train a secondary reward model focused on safety from binary action critiques. This is in addition to the main reward model, which is trained from comparisons in their approach. Note that this secondary safety model could, in principle, be trained with any of the feedback types discussed here, using methods identical to the ones used for reward learning.
3.2.2 Comparisons
@@
Binary comparisons and rankings are among the most common types of feedback. The defining features of comparisons are that the human passively observes the behavior (involvement), gives relative feedback on multiple instances (arity), and does so explicitly (explicitness) with an evaluative intent. It is most commonly given on a segment (granularity), but other granularities are possible in principle [60]. Similarly, comparisons are commonly requested on an instance level (abstraction), but this is not a requirement.
@@
Comparisons were first used for direct policy learning [3, 53], but were later extended to the reward-learning setting [273, 58]. The most common setting [58] relies on pairwise comparisons of trajectory segments, but comparisons of individual states or actions were also considered in early PbRL works [83].
@@
This basic setting has been extended and modified in various ways. To reduce noise in the labels, it is common to extend the binary choice by giving the labelers the option to avoid the hard choice and instead indicate incomparability, uncertainty, or perceived similarity [117]. The most common interpretation of this intermediate option is “equally preferable” [169], e.g., as if each trajectory had an equal probability of being preferred in the preference predictor. [119] instead provide two intermediate options: The previously discussed “equally preferable” and an “incomparable” option which results in simply omitting the query from the data set. In contrast to this, [257] explicitly state that they do not allow for the equally preferred option, arguing that these cases are rare enough not to matter very much. Another line of research suggests that more precision in the expression of pairwise preferences, such as softening the hard binary choice to scalar feedback indicating the strength of a preference [266], can be beneficial for preference learning. Other extensions change the pairwise setting to choices among larger choice sets [296] or even full rankings [188, 39, 200]. [296] note that for language tasks, a larger choice set can amortize the time needed for a labeler to get acquainted with the context necessary to understand a query. [18] propose to use hierarchical queries, i.e., a sequence of pairwise comparisons that build up on each other.
@@
Pairwise comparisons provide many benefits over absolute ratings, such as reduced bias, inconsistencies, and subjectivity [283], but also convey relatively little information per label. [250] study the weaknesses of pairwise-comparison-based reward learning. Since the amount of information provided for each label is small, these models are prone to causal confusion, i.e., misattributing reward to noise and misidentifying the reward function.
3.2.3 Inter-Temporal Feedback
@@
Reward sketching, as introduced by [42], involves users sketching a visual representation of the reward function over time. This type of feedback, which can be given by sketching a graph with the mouse while watching a behavior recording, provides intuitive, per-timestep reward annotations. [214] also adopted this approach, referring to it as “one of the highest-bandwidth feedback mechanisms currently available”.
Inter-temporal preferences were introduced by [2]. In this setting, humans give feedback on multiple points of a trajectory, indicating whether an agent makes progress towards or regresses from a goal. This is then interpreted as preferences relative to the other labeled and unlabelled points. The authors note that one potential downside of this feedback type is that labelers may tend to give preferences on short-term actions that are easy to judge, failing to communicate long-horizon preferences. [61] propose a similar type of feedback, in which humans segment a trajectory into good and bad parts. This makes it possible to derive many state-action labels from few segmentation points.
3.2.4 Task Descriptions
Task descriptions consist of direct, concept-level descriptions of the task the agent is supposed to complete. This is most commonly conveyed in the form of information about the reward function, i.e., by proxy rewards or reward queries. Descriptive feedback does not generally refer to any particular behavior instance but instead gives global direction for the entire task. However, in line with our selection criteria (Section 1.3), we only consider task descriptions that are interactive and online within the context of one or multiple observations to fill holes in the initial description. The defining features of task descriptions are that the labeler passively observes the agent’s behavior (involvement) and gives feedback explicitly (explicitness) on a feature-level (abstraction) with descriptive intent (intent). Descriptive feedback may be given with respect to a single or multiple instances (arity), although it generally refers to multiple instances. It is most commonly given on an episode (granularity), but other granularities are possible in principle.
Task descriptions are most commonly provided in the form of incomplete reward functions. [112] demonstrate this using proxy reward functions, which are preliminary reward functions that might not cover all edge cases, to guide the agent toward learning the actual reward function. Alternatively, [185, 108] suggest querying about the reward function. They allow users to choose from a set of understandable, linear proxy rewards or to specify which features are more critical in the linear reward structure.
3.2.5 Social Behavior
@@
[62] propose a framework to learn reward functions from such social behavior. They suggest a two-phase training setup. In the first phase, they ground the implicit feedback by use of incentives, i.e.,  they incentivize humans to have a known objective. After learning a mapping from feedback to reward, they use regular RL techniques to learn a policy. Note that the learned reward function can be seen as conditional on human implicit feedback and, therefore, they require a human in the loop throughout training.
3.2.6 Improvements
@@
Interventions can be considered to be an instance of the shared autonomy setting since the agent and the user share autonomy to reach a common goal. This setting is studied by [2], who ask humans to intercede on agent failure. They use this to collect targeted demonstrations where the agent is the weakest and to identify challenging situations for their evaluations. The gathered data is then used for behavior cloning and reward model training. The fact that a correction occurred is not directly used as feedback. In contrast to this [171] observe that the interventions themselves are intentional and therefore carry information about the true objective. [172] propose to incorporate uncertainty for active learning and risk-sensitive deployment in this setting. [160] further extend the intervention setting to learn from a sequence of correlated physical corrections without needing to wait until the trajectory is completed.
@@
The correction case is closely related to the setting of coactive learning [234], in which a learning system repeatedly proposes a solution which a user may correct to reach a common goal. [122] treat corrections as a demonstration while [125] propose (but do not evaluate) an alternative interpretation of inferring implicit preferences from comparisons by assuming the corrected trajectory is preferred over the original one.
3.2.7 E-Stops
@@
Emergency stops (e-stops) [91] are an active type of feedback. In this type of feedback, the human may intervene with the agent’s behavior by stopping it, i.e., they may choose to stop the agent’s current trajectory at any point. This is closely related to interventions but, in contrast to those, does not suggest an alternative action. The defining features of e-stops are that the human both passively observes the agent’s behavior (involvement), gives absolute feedback on a single instance (arity) on the instance level (abstraction) and does so implicitly (explicitness) as a side-effect of regular interaction. The intent is literal due to the implicit nature. For the purposes of intervention, the human usually observes the full episode (granularity). Due to the small amount of infrequent information they provide, e-stops should only be considered as a supplementary feedback type.
@@
E-stops are primarily intended to prevent bad behavior and only implicitly convey information about the correct behavior. This interaction and the arising incentives have been formalized in the form of the “off-switch game” by [107]. [125] propose to interpret this as a form of reward-rational feedback, where the ‘off’ choice maps to the trajectory with the robot remaining still after the off switch has been triggered. [131] demonstrate that a robot can learn to navigate using such feedback.
3.2.8 Importance
@@
One way to convey this information is by labeling salient parts of a visual input. This is explored by [101], who augment pairwise comparisons with manually annotated visual saliency maps, informing the algorithm which parts of the visual input contributed to the decision. They leverage these annotations for data augmentation by assuming that random perturbations to irrelevant (non-salient) regions do not impact the human preferences. [19] take an even more direct approach by combining comparative feedback with direct feature queries, i.e., asking the user which feature is important for inferring the reward.
3.2.9 Feature Traces

### v2

@@ 3 Feedback
[…]
Figure 2: Highlighting the components of the RLHF loop discussed in this section.
@@
A feedback type is a kind of interaction in which a human conveys some information about their preferences. Examples include pairwise comparisons and direct critiques. This section is concerned with the many ways in which human feedback can be expressed and used. Several previous works have already studied and sorted feedback types by listing the most common ones (Jeon et al., 2020; Yuan et al., 2024) and discussing their attributes and dimensions (Metz et al., 2023; Lindner & El-Assady, 2022). The attributes and classes described in this section build upon this prior work and can be considered a synthesis and extension of it.
@@ 3.1 Attributes of Feedback Types
Feedback types may differ on many dimensions, some of which relate to the way feedback is given (arity, involvement), others to the form of the query instance it is given on (granularity, abstraction), and yet others to features of the human interaction (intent, explicitness). The attributes we discuss in this section are based on the framework proposed by Metz et al. (2023). We have adjusted and added terminology where it aids clarity, generalized the distinction between relative and absolute feedback to ‘arity’, and added the categories of co-generative involvement and literal intent. Furthermore, we systematically analyze a set of exemplary feedback types in the next section, expanding on the initial examination of a smaller range of abstract classes by Metz et al. (2023). In the following, we will introduce each of the six attributes in more detail.
@@
Feedback may also differ on the granularity of the instances being evaluated. This ranges from whole episode recordings over partial segments to feedback on individual steps (i.e., states, actions, or state-action pairs). A more coarse-grained granularity has the advantage of giving humans more context and getting feedback for larger sections of behavior but also poses credit assignment problems. Finer-grained feedback is much easier to learn from and simplifies credit assignment, but is often impractical or tedious for humans to provide. It is also possible to strike a compromise, as demonstrated by Guan et al. (2021) who propose to use queries on step granularity, but batch queries within an episode to provide additional context and make it easier to give feedback. Note that we only classify a type of feedback as “episode” granularity if it requires entire episodes. If it is compatible with partial segments as well, we classify it as “segment” even if the discussed source paper uses entire episodes.
@@
Intent
The assumed human intent can be important for feedback processing. A human may be evaluative, instructive, or descriptive in their explicit feedback, while they are generally literal in their implicit feedback. Evaluative, instructive, and descriptive feedback is pedagogical in nature, aiming to teach a reward function, whereas literal feedback is a byproduct of a human actor’s efforts to optimize the reward function directly. Descriptive feedback is often given on the level of the task, e.g., through a partial reward function. The other types of intent, in contrast, are generally given within the context of a particular instance of behavior.
The distinction between literal and pedagogical feedback was introduced by Milli & Dragan (2020). They argue that humans with pedagogical intent (i.e., evaluative, instructive, or descriptive) may act differently compared to humans with literal intent. Even though they find that assuming the (wrong) literal intent can still lead to better reward inference, it still indicates that it can be important to know this intent to choose the right human model (see Section 5.1.5).
@@ 3.2 Common Classes
Table 4: An overview of the common classes and their defining attributes. The feedback classes are grouped into primary classes, which can be used on their own to learn a reward model, supplementary (sup.) classes, which can be used in conjunction with a primary class, and representation-focused (rep.) classes, which can be used to learn a representation prior to learning a reward model. When is specified for an attribute, this indicates that it is not a defining feature of the class and may vary in different instantiations.
Class Granularity Involvement Arity Abstr. Intent Expl.
[…]
Primary
@@
Proxy Rewards Episode Observed Feature Descriptive Explicit
@@
Natural Language Observed Unary Descriptive Explicit
[…]
Sup.
@@
3.2.1 Primary Feedback Classes
This section introduces common feedback classes which can be used on their own to learn a reward model. The classes are critique, comparisons, inter-temporal feedback, proxy rewards, social behavior, improvements, and natural language.
Critique
@@ 3.2.1 Primary Feedback Classes
There are many examples of critique feedback in the literature. Xiao et al. (2020) employ binary feedback on individual state and action pairs. Although they learn a shaping reward that complements an environment reward signal, the same technique could be used without environment reward. Huang et al. (2023) extend this to multi-label feedback, allowing the user to distinguish between a regular good or bad action and a terminal action that achieves the goal or fails to do so. They map these classes to scalar values and then learn a reward model by regression. Wang et al. (2020) present an approach to learning a reward model from noisy critiques in the form of human physiological signals (brain signals) using active querying. In contrast to this action-level feedback, Fu et al. (2018b); Singh et al. (2019) rely on binary outcome success labels. Fu et al. (2018b) introduce the basic approach, which Singh et al. (2019) extend by moving to an off-policy setting and including online queries, thereby reducing the reliance on many positive examples by interactively correcting false positives.
@@
In addition to learning the main reward function, critique can also be used for safety evaluation. Cosner et al. (2022) train a secondary reward model focused on safety from binary action critiques. This is in addition to the main reward model, which is trained from comparisons in their approach. Note that this secondary safety model could, in principle, be trained with any of the feedback types discussed here, using methods identical to the ones used for reward learning.
Comparisons
@@
Binary comparisons and rankings are among the most common types of feedback. The defining features of comparisons are that the human passively observes the behavior (involvement), gives relative feedback on multiple instances (arity), and does so explicitly (explicitness) with an evaluative intent. It is most commonly given on a segment (granularity), but other granularities are possible in principle (Cosner et al., 2022). Similarly, comparisons are commonly requested on an instance level (abstraction), but this is not a requirement.
@@
Comparisons were first used for direct policy learning (Akrour et al., 2011; Cheng et al., 2011), but were later extended to the reward-learning setting (Wirth et al., 2016; Christiano et al., 2017). The most common setting (Christiano et al., 2017) relies on pairwise comparisons of trajectory segments, but comparisons of individual states or actions were also considered in early PbRL works (Fürnkranz et al., 2012) and comparisons can even be extended to more abstract trajectory features (Pinsler et al., 2018).
@@
This basic setting has been extended and modified in various ways. To reduce noise in the labels, it is common to extend the binary choice by giving the labelers the option to avoid the hard choice and instead indicate incomparability, uncertainty, or perceived similarity (Holladay et al., 2016). This option is commonly interpreted as “equally preferable”, e.g., as if each trajectory had an equal probability of being preferred in the preference predictor. It is alco common, however, to additionally provide an “incomparable” option which results in simply omitting the query from the data set (Christiano et al., 2017; Ibarz et al., 2018). In contrast to this, Verma et al. (2023) explicitly state that they do not allow for the equally preferred option, arguing that these cases are rare enough not to matter very much. Another line of research suggests that more precision in the expression of pairwise preferences, such as softening the hard binary choice to scalar feedback indicating the strength of a preference (Wilde et al., 2021), can be beneficial for preference learning. Other extensions change the pairwise setting to choices among larger choice sets (Ziegler et al., 2020) or even full rankings (Myers et al., 2021; Brown et al., 2020; Ouyang et al., 2022). Jain et al. (2015) compare different kinds of re-ranking feedback (select first which is better than top, choose best from top-5, choose best from random set). Their study suggests that the first two are roughly equivalent, while the latter is much less informative. Askell et al. (2021) evaluate binary comparisons as well as ranking, but finds that ranking performs better. Ziegler et al. (2020) note that for language tasks, a larger choice set can amortize the time needed for a labeler to get acquainted with the context necessary to understand a query. Basu et al. (2019) propose to use hierarchical queries, i.e., a sequence of pairwise comparisons that build up on each other.
@@
Pairwise comparisons are generally easier to supply than ratings, even when the feedback is given by a foundation model (Wang et al., 2024b). While absolute feedback is generally dependent on a policy serving as an implicit baseline (MacGlashan et al., 2017), the explicit baseline in pairwise comparisons can be more easily controlled. Therefore, comparisons provide many benefits over absolute ratings, such as reduced bias, inconsistencies, and subjectivity (Yannakakis & Martínez, 2015). On the flip-side, comparisons convey relatively little information per label. Tien et al. (2023) study the weaknesses of pairwise-comparison-based reward learning. Since the amount of information provided for each label is small, these models are prone to causal confusion, i.e., misattributing reward to noise and misidentifying the reward function.
Inter-Temporal Feedback
@@
Reward sketching, as introduced by Cabi et al. (2020), involves users sketching a visual representation of the reward function over time. This type of feedback, which can be given by sketching a graph with the mouse while watching a behavior recording, provides intuitive, per-timestep reward annotations. Rahtz et al. (2022) also adopted this approach, referring to it as “one of the highest-bandwidth feedback mechanisms currently available”.
Inter-temporal preferences were introduced by Abramson et al. (2022). In this setting, humans give feedback on multiple points of a trajectory, indicating whether an agent makes progress towards or regresses from a goal. This is then interpreted as preferences relative to the other labeled and unlabelled points. The authors note that one potential downside of this feedback type is that labelers may tend to give preferences on short-term actions that are easy to judge, failing to communicate long-horizon preferences. Cui & Niekum (2018) propose a similar type of feedback, in which humans segment a trajectory into good and bad parts. This makes it possible to derive many state-action labels from a few segmentation points.
Proxy Rewards
Proxy rewards are partial or inaccurate reward functions that convey information about the task the agent is supposed to complete but may not induce optimal behavior. This form of feedback does not generally refer to any particular behavior instance but instead gives global direction for the entire task. However, in line with our selection criteria (Section 1.3), we only consider proxy reward feedback that is interactive and online within the context of one or multiple observations to fill holes in the initial description. The defining features of proxy reward feedback is that the labeler passively observes the agent’s behavior (involvement) and gives feedback explicitly (explicitness) on a feature-level (abstraction) with descriptive intent (intent). Proxy reward feedback may be given with respect to a single or multiple instances (arity), although it generally refers to multiple instances. It is most commonly given on an episode (granularity), but other granularities are possible in principle.
The work by He & Dragan (2021) exemplifies this form of feedback by using proxy reward functions, i.e., preliminary reward functions that might not cover all edge cases, to guide the agent toward learning the actual reward function. Alternatively, Mindermann et al. (2018); Hadfield-Menell et al. (2017b) suggest querying about the reward function. They allow users to choose from a set of understandable, linear proxy rewards or to specify which features are more critical in the linear reward structure. In a related setting, Guan et al. (2023) lets the user specify changes to a current symbolic reward function in the form of changes to target attributes (e.g., walking speed).
Social Behavior
@@
Cui et al. (2021) propose a framework to learn reward functions from such social behavior. They suggest a two-phase training setup. In the first phase, they ground the implicit feedback by use of incentives, i.e.,  they incentivize humans to have a known objective. After learning a mapping from feedback to reward, they use regular RL techniques to learn a policy. Note that the learned reward function can be seen as conditional on human implicit feedback and, therefore, they require a human in the loop throughout training.
Improvements
@@
Interventions can be considered to be an instance of the shared autonomy setting since the agent and the user share autonomy to reach a common goal. There are two main ways to leverage interventions: One is to learn form the occurrence of an intervention itself that the prior behavior was suboptimal, as done, e.g., by Luo et al. (2024) by directly inferring negative rewards from interventions without learning a reward model, and the other is to learn from the corrected trajectory. The latter setting is studied by Abramson et al. (2022), who ask humans to intercede on agent failure. They use this to collect targeted demonstrations where the agent is the weakest and to identify challenging situations for their evaluations. The gathered data is then used for behavior cloning and reward model training. The fact that a correction occurred is not directly used as feedback. In addition to the above distinction, interventions can come in two main formats: Either by overtaking control from the agent to correct its behavior or by physically correcting the agent’s movements. Losey et al. (2022) proposes to learn a reward model from such physical corrections, a method that was later extended (Losey & O’Malley, 2018) to incorporate uncertainty for active learning and risk-sensitive deployment. Li et al. (2021) further extend this setting to learn from a sequence of correlated physical corrections without needing to wait until the trajectory is completed.
@@
The correction case is closely related to the setting of coactive learning (Shivaswamy & Joachims, 2015), in which a learning system repeatedly proposes a solution which a user may correct to reach a common goal. Jain et al. (2015) treat corrections as a demonstration while Jeon et al. (2020) propose (but do not evaluate) an alternative interpretation of inferring implicit preferences from comparisons by assuming the corrected trajectory is preferred over the original one.
Natural Language
Natural language is a versatile form of feedback that can be used to convey a wide range of information. It cannot only be used to express preferences, but also to suggest concrete changes. Natural language feedback may be given on any granularity, at any level of abstraction. Its defining features are that it is given explicitly (explicitness) in the context of a single (arity) observed behavior or policy (involvement) with descriptive intent.
While much work in language for RL focuses on the problem of learning a policy that follows natural-language instructions in a non-interactive setting (Hermann et al., 2017; Nair et al., 2021), we focus on works where language is used as feedback, not for task definition. Note, however, that RLHF can be a useful tool for learning a language-conditioned policy, regardless of whether or not the feedback itself is in the form of natural language. As an example of this, Abramson et al. (2022) use RLHF in an interactive agents setting, where one agent (responsible for solving problems) is controlled by the learned policy and another agent (responsible for setting tasks) is controlled by the human.
Language may be interpreted as a form of feedback in the reward-rational choice setting (Jeon et al., 2020), where an utterance implies a set of trajectories compatible with the utterance (grounding) and the human can choose an utterance that maximizes the probability of a desired action (i.e., the human is assumed to be specific which leads to pragmatic reasoning). This can be used to infer a reward which may have caused the human to provide the utterances they did provide. A similar approach is followed, e.g., by (Sumers et al., 2022b).
An alternative way of incorporating language feedback is by extracting more structured forms of feedback from the language. An example is to use sentiment analysis to extract information about the reward function from language feedback, as proposed by Sumers et al. (2021).
Language feedback can also be interpreted by an LLM directly, without learning a reward model. This is demonstrated in the concurrent works by Ma et al. (2024) and Xie et al. (2024) who propose to learn a symbolic reward model in the form of a language-model written piece of (Python) code. They use natural language feedback to improve the reward model based on observations of the agent’s behavior induced by the previous version of the reward model.
In addition to the way language feedback is used, the way it is elicited can also have an impact on the learning process. Sumers et al. (2022a) study the relative merits of natural language instructions and descriptions of the desired outcome. They find that instructions tend to work better for low-autonomy settings, while descriptions are more effective in high-autonomy settings.
3.2.2 Supplementary Classes
Merge this with representation-specific, highlight distinction direct feedback about preferences vs feedback about representation in section introduction
This section introduces two feedback classes, e-stops and importance, that can be used in conjunction with a primary class to learn a reward model. These classes are not sufficient to learn a reward model on their own, but can supplement a primary feedback type.
E-Stops
@@ 3.2.2 Supplementary Classes
Emergency stops (e-stops) (Ghosal et al., 2023) are an active type of feedback. In this type of feedback, the human may intervene with the agent’s behavior by stopping it, i.e., they may choose to stop the agent’s current trajectory at any point. This is closely related to interventions but, in contrast to those, e-stops do not suggest an alternative action. The defining features of e-stops are that the human both passively observes the agent’s behavior (involvement), gives absolute feedback on a single instance (arity) on the instance level (abstraction) and does so implicitly (explicitness) as a side-effect of regular interaction. The intent is literal due to the implicit nature. For the purposes of intervention, the human usually observes the full episode (granularity). Due to the small amount of infrequent information they provide, e-stops should only be considered as a supplementary feedback type.
@@
E-stops are primarily intended to prevent bad behavior and only implicitly convey information about the correct behavior. This interaction and the arising incentives have been formalized in the form of the “off-switch game” by Hadfield-Menell et al. (2017a). Jeon et al. (2020) propose to interpret this as a form of reward-rational feedback, where the ‘off’ choice maps to the trajectory with the robot remaining still after the off switch has been triggered (see Section 5.1.1). Kahn et al. (2021) demonstrate that a robot can learn to navigate using such feedback.
Importance
@@
One way to convey this information is by labeling salient parts of a visual input. This is explored by Guan et al. (2021), who augment pairwise comparisons with manually annotated visual saliency maps, informing the algorithm which parts of the visual input contributed to the decision. They leverage these annotations for data augmentation by assuming that random perturbations to irrelevant (non-salient) regions do not impact the human preferences. Basu et al. (2018) take an even more direct approach by combining comparative feedback with direct feature queries, i.e., asking the user which feature is important for inferring the reward.
3.2.3 Representation-Specific Classes
While the previous classes of feedback types are all aimed at directly learning a reward function, there are also classes of feedback types that do not directly learn a reward function but rather help to learn a better representation.
Feature Traces